# Negative Dependence as a toolbox for machine learning : review and new developments

### v1

@@ 8.4 Applications in signal analysis
The connection to the zeros of GAFs provides a useful approach for studying the stochastic geometry of spectrogram level sets of a signal. In [GLS22], Ghosh et al. provided a theoretical guarantee for using GAFs in the problem of signal detection. To elaborate, let $f$ be a noisy signal generated as
@@
$\mathcal{L}(a)$ ::=:= $\{z\in\mathbb{C}:|\mathcal{F}_{\phi}f(z)|\geq a\},\quad a\geq 0$ (8.6)
$\mathcal{L}_{L}(a)$ ::=:= $\mathcal{L}(a)\cap\mathbb{B}_{L},\quad\text{where }\mathbb{B}_{L}:=\{z\in\mathbb{C}:|\mathrm{Re}(z)|\leq L,|\mathrm{Im}(z)|\leq L\}.$ (8.7)
Utilising Equation (8.3) and techniques from maxima of Gaussian random fields, Ghosh et al. in [GLS22] showed the following result for the spectrogram level set of the white noise
@@
$\mathbb{P}\Big{(}\sup_{z\in\mathbb{B}_{L}}|\mathcal{F}_{\phi}\xi(z)|\leq\sqrt{2}(14K+\tau)\sqrt{\log L}\Big{)}\geq 1-4\exp\Big(-\frac{\tau^{2}}{2\pi}\log L\Big{missing})$
@@
and the supremum is attained at $|z|=\sqrt{k/\pi}$ (see [GLS22]). Combining the ingredients above, we have the following theorem.
@@
Theorem 8.10 provides useful technical tools in signal detection. In particular, the authors in [GLS22] present two applications: one in hypothesis testing and one in signal estimation. For the first application, they design an effective hypothesis test to distinguish between pure white noise and the presence of a fundamental signal. They also provide theoretical guarantees for its effectiveness using non-asymptotic bounds on the power of the test. For the second application, they demonstrate an estimation procedure for identifying a fundamental signal, if present, and provide error bounds on the accuracy of this estimation based on the sample size. For more details, we refer readers to Section 4 of [GLS22].
@@ 9 DPPs for feature selection
This section is based on [BBC20]. It is well known that the best (e.g. in Frobenius norm) approximation of a given matrix $X\in\mathbb{R}^{N\times d}$ by a matrix of smaller rank $k\ll d$ is obtained by principal component analysis (PCA). The main drawback from this is that one loses the explainability of the features (columns), which can be particularly troublesome in some areas. Column subset selection aims at resolving this particular issue: for the matrix $X$ containing $d$ features about $N$ individuals, feature selection consists in selecting $k\ll d$ columns that are the most relevant. Finding the k-columns which contributes the most to the matrix is a complex task, and randomized methods have proven to be particularly efficient in this regard.
@@
Solving the problem of finding a good candidate $S$ is a challenge. Deterministic algorithms exist such as rank revealing QR (RRQR) decomposition, see [BMD09], which have their own caveats. Here we choose to focus on non-deterministic methods.
@@ 9.1.1 Length squared importance sampling
The first idea to randomly select columns is to give them importance scores as their norm. This algorithm was introduced by Drineas et al. in [DFK+04]. More precisely, the idea is to sample $s$ indices, from a multinomial distribution on $[d]$, where the probability to sample $i$ is $p_{i}\propto\lVert X_{:,i}\rVert^{2}$.
Theorem 3 in [DFK+04] states that:
@@ 9.1.2 Leverage score importance sampling
Another importance score that can be considered is the leverage score. This method was introduced by Drineas et al. in [DMM08]. The $k$-leverage scores distribution is the multinomial distribution on $[d]$ with parameters $p_{i}=l_{i}^{k}/k$. It is easy to see that the leverage scores sum to $k$ so that this definition gives a proper probability distribution. Besides, for $\mathcal{P}_{k}=\operatorname{Span}(V_{k})$, when $(e_{i})_{i\in[d]}$ is the canonical basis of $\mathbb{R}^{d}$, and $\theta_{i}$ is the angle between $\mathcal{P}_{k}$ and $e_{i}$, some linear algebra manipulations show that $l_{i}^{k}=\cos^{2}(\theta_{i})$. So selecting the columns according to the leverage score sampling is in fact choosing the coordinates $e_{i}$ that are the most aligned with $\mathcal{P}_{k}$. PCA selects exactly $\mathcal{P}_{k}$, hence, selecting a subspace which is almost aligned with $\mathcal{P}_{k}$ should yield a good approximation.
Theorem 3 in [DMM08] states that, for $\epsilon>0$, $\delta>0$, if the number of sampled columns $s$ satisfies $s\geq 4000(k^{2}/\epsilon^{2})\log(1/\delta)$, then under the $k$-leverage score sampling it holds:
@@ 9.2.1 Volume sampling
One way to introduce correlation between the columns of $X$ is through volume sampling. The idea is here to use the kernel $X^{T}X$ and sample a $k$-DPP from it. This method will be called volume sampling (VS) in the following. Notice that the diagonal terms of the kernel are exactly the squared lenghts of the columns of $X$. Volume sampling was first introduced by Deshpande et al. in [DRVW06] and they established the following result:
@@
[[DRVW06] Theorem 1.3] Assume $\mathcal{S}$ is sampled according to the volume sampling method. Then:
@@
Recently, [DKM21] have worked on these bounds and they proved the following result. Notice that the bound is sharp in the worst case scenario, so they play on the decay of the singular values to achieve a better bound in some cases.
Proposition 9.2 ([DKM21] Theorem 1).
@@
Let $20<\epsilon\leq 1/$ 2, $s<r$ and $t_{s}=s+sr_{s}(X)$, then for $s+\frac{7}{\epsilon^{2}}\ln(\frac{1}{\epsilon^{2}})\leq k\leq t_{s}-1$, the following holds:
@@
This upper bound is rather complicated to interpret in full generality. In their paper [DKM21], the authors show that it actually improves 9.1 in cases where the singular values decay at a polynomial or an exponential rate.
@@ 9.2.2 DPP sampling
The last algorithm we wanted to mention is taken from [BBC20]. Their idea is to force the alignment between the subspace generated by the $k$ selected coordinates and the subspace generated by simultaneously force each principle angles to be the smallest possible. In this regard, consider $S\subset[d]$, $\lvert S\rvert=k$ and denote $\mathcal{P}_{S}=\operatorname{Span}\{e_{j}|j\in S\}$. Let $\mathbb{\theta}=(\theta_{i})_{1\leq i\leq k}$ be the principal angle between $\mathcal{P}_{k}$ and $\mathcal{P}_{S}$. Then it can be shown that: $\prod\limits_{i=1}^{k}\cos(\theta_{i})^{2}=\mathrm{Det}(V_{S,[k]}V_{S,[k]}^{T})$. Hence sampling from an elementary DPP (of rank $k$) with kernel $K=V_{k}V_{k}^{T}$ emerges naturally from the geometry of the problem. This is what will be called DPP sampling in the following.
@@
Proposition 9.3 ([BBC20] Proposition 16).
@@
Proposition 9.4 ([BBC20] Proposition 17).
@@
The proof scheme for this result is quite similar to the one of 9.1 with added technicalities. More precisely, notice that the inequality which was used there is rather loose, and a better understanding of the decay of the singular values of $X$ would give a better bound. Besides, this projection DPP has a strong connection to the principal angles as mentioned previously, and in a certain way, sampling from this DPP makes us find the columns that are best aligned with the principal eigenspaces. Using these facts, and correctly manipulating the symmetric polynomials lead to this result. We refer the reader to [BBC20] for more details.
@@
It is worth mentioning that it is possible to reduce the sparsity by looking at some ”effective sparsity”, and conditioning on sampling indices for which the leverage score contributes significantly to the whole. All of these results can also be extended when working with the spectral norm instead of the Frobenius one, see [BBC20].
@@ 10.1 Background
Pruning neural networks consists in reducing the number of neurons in the network. Neural network pruning is a central question when it comes to optimizing neural network performances. Mariet and Sra developed a method, DIVNET in [MS17], for pruning neural network using DPP which has proven to be empirically efficient and is independent of the choice of other parameters such as the activation function or the number of hidden layers. We will describe the method for feed forward neural networks.
@@
Looking at it through the lens of statistical mechanics offers some insight. This framework was already introduced in [ACZ+21]. We aim at offering a comprehensive overview of the methods coming from statistical mechanics applied to the study of neural network pruning. We also extend and formalize the results from [ACZ+21] to more general settings.
The analysis presented in this section was originally developed in the masters thesis [Pet23] of co-author V. Petrovic .
@@ 10.2 The student/teacher framework
The student/teacher framework is a common setting in statistical mechanical analysis of neural networks. The full theory behind this model is fully explained and detailed in [EVdB01]. The model that we will use in this section comes from [GAS+20] in the noiseless regime.
The framework is the following, assume we are given a data set $(X_{\mu})_{1\leq\mu\leq p}$ of i.i.d. Gaussian variables which take values in $\mathbb{R}^{d}$, of mean 00 and variance matrix $I_{d}$. Let us get two neural networks respectively called the teacher and the student. The teacher network acts as a black box, to which we have access to the activation function and the number of layers and neurons per layer. It gets as inputs the data set $(X_{\mu})_{1\leq\mu\leq p}$ and assigns them outputs $(Y_{\mu})_{1\leq\mu\leq p}$, where if $1\leq\mu\leq p$, $Y_{\mu}=\phi^{*}(X_{\mu})$, where $\phi^{*}$ is a deterministic function. The student network is then another neural network, on which we have full control of the architecture, that we train to reproduce the results of the teacher network. In other words, the student network performs supervised learning with inputs $(X_{\mu},Y_{\mu})_{1\leq\mu\leq p}$. Finally, let $g$ be the activation function, which is chosen to be the sigmoid function $\operatorname{erf}$. For the rest of the section, we will use $i,j,k$ to describe nodes from the student network and $n,m$ to describe nodes from the teacher network.
@@
$\epsilon_{g}(\phi)=\dfrac{1}{2}\innerproduct{(\phi(X)-\phi^{*}(X))^{2}}{(\phi(X)-\phi^{*}(X))^{2}}$
where $\innerproduct{\cdot}{\cdot}$ is the average over the input distribution.
@@
The whole analysis will revolve around the dynamics established by Goldt et al. in [GAS+20] .
@@ 10.3 Dynamics of the macroparameters in online learning
Goldt et al establish very general results on the evolution of the paramaters in online learning. In the following, we will focus on a specific case mentioned by the authors of [GAS+19] and further assume that:
@@
$\forall 1\leq i,k\leq K\text{, }I_{2}(i,k)=\innerproduct{g(\lambda_{i})g(\lambda_{k})}{g(\lambda_{i})g(\lambda_{k})}=\dfrac{1}{\pi}\arcsin{\dfrac{Q_{ik}}{\sqrt{1+Q_{ii}}\sqrt{1+Q_{kk}}}}$ )
$\forall 1\leq i\leq K\text{, }\forall 1\leq n\leq M\text{, }I_{2}(i,n)=\innerproduct{g(\lambda_{i})g(\rho_{n})}{g(\lambda_{i})g(\rho_{n})}=\dfrac{1}{\pi}\arcsin{\dfrac{R_{in}}{\sqrt{1+Q_{ii}}\sqrt{1+T_{nn}}}}$ )
@@
$\forall 1\leq m,n\leq M\text{, }I_{2}(m,n)=\innerproduct{g(\rho_{m})g(\rho_{n})}{g(\rho_{m})g(\rho_{n})}=\dfrac{1}{\pi}\arcsin{\dfrac{T_{mn}}{\sqrt{1+T_{nn}}\sqrt{1+T_{mm}}}}$ )
The authors of [GAS+20] establish the following theorem:
@@ 10.4 Application to neural network pruning
We assume that $T_{nm}=\delta_{nm}$ (as [ACZ+21]), i.e., that we work in the ideal case where the weights in the teacher network are a subfamily of orthonormal vectors. Assume further that both hypothesis (H1) and (H2) from section 10.3 hold.
@@
This is a generalization of [ACZ+21] in the following sense:
@@ 10.5.3 DPP sampling
Equivalently, if $i,j\in G_{m}$, $i\neq j$, then ($\mathbb{P}(\{i\in\mathcal{S}\}\cap\{j\in\mathcal{S}\})=0$.
@@
  • the first example is by taking $L_{ij}=Q_{ij}$. Notice that, since $Q_{ij}=\frac{1}{N}w_{i}^{T}w_{j}$, $Q$ is a Gram-matrix so it is semi definite positive, and is an admissible choice for $L$.
@@ 11 A quantum sampler for DPPs
While her thesis initially was concerned with finding a stochastic description of this bunching effect, without the semiclassical description of the electric field as a function-valued stochastic process, Odile Macchi then turned to model a similar situation where photons are replaced by fermions, like electrons. There, she found out that detection events are negatively correlated, evidencing an anti-bunching effect that was later experimentally observed. Macchi formalized DPPs precisely to model that detection process [Mac72]; see also [BFB+22] for a modern review. Mathematically, DPPs arise from the combined action of Born’s rule –the quantum physics axiom that lab measurements are draws from a random variable parametrized by two operators on a Hilbert space– and the canonical anti-commutation relations satisfied by the building blocks of operators describing fermions.
In this section, we walk in Macchi’s footprints and illustrate how DPPs are related to quantum fermionic models in a simple setting, showing in passing that we can design an algorithm to sample a given DPP on a quantum computer. This section paraphrases [BFF24], to which we refer the reader for more details and important references, such as [JSK+18, KP22].

### v4

@@ 8.4 Applications in signal analysis
The connection to the zeros of GAFs provides a useful approach for studying the stochastic geometry of spectrogram level sets of a signal. In [GhoshGAF], Ghosh et al. provided a theoretical guarantee for using GAFs in the problem of signal detection. To elaborate, let $f$ be a noisy signal generated as
@@
$\mathcal{L}(a)$ $:=$ $\{z\in\mathbb{C}:|\mathcal{F}_{\phi}f(z)|\geq a\},\quad a\geq 0$ (8.6)
$\mathcal{L}_{L}(a)$ $:=$ $\mathcal{L}(a)\cap\mathbb{B}_{L},\quad\text{where }\mathbb{B}_{L}:=\{z\in\mathbb{C}:|\mathrm{Re}(z)|\leq L,|\mathrm{Im}(z)|\leq L\}.$ (8.7)
Utilising Equation (8.3) and techniques from maxima of Gaussian random fields, Ghosh et al. in [GhoshGAF] showed the following result for the spectrogram level set of the white noise
@@
[…]
@@
and the supremum is attained at $|z|=\sqrt{k/\pi}$ (see [GhoshGAF]). Combining the ingredients above, we have the following theorem.
@@
Theorem 8.10 provides useful technical tools in signal detection. In particular, the authors in [GhoshGAF] present two applications: one in hypothesis testing and one in signal estimation. For the first application, they design an effective hypothesis test to distinguish between pure white noise and the presence of a fundamental signal. They also provide theoretical guarantees for its effectiveness using non-asymptotic bounds on the power of the test. For the second application, they demonstrate an estimation procedure for identifying a fundamental signal, if present, and provide error bounds on the accuracy of this estimation based on the sample size. For more details, we refer readers to Section 4 of [GhoshGAF].
@@ 9 DPPs for feature selection
This section is based on [belhadji_determinantal_2020]. It is well known that the best (e.g. in Frobenius norm) approximation of a given matrix $X\in\mathbb{R}^{N\times d}$ by a matrix of smaller rank $k\ll d$ is obtained by principal component analysis (PCA). The main drawback from this is that one loses the explainability of the features (columns), which can be particularly troublesome in some areas. Column subset selection aims at resolving this particular issue: for the matrix $X$ containing $d$ features about $N$ individuals, feature selection consists in selecting $k\ll d$ columns that are the most relevant. Finding the k-columns which contributes the most to the matrix is a complex task, and randomized methods have proven to be particularly efficient in this regard.
@@
Solving the problem of finding a good candidate $S$ is a challenge. Deterministic algorithms exist such as rank revealing QR (RRQR) decomposition, see [boutsidis_improved_nodate], which have their own caveats. Here we choose to focus on non-deterministic methods.
@@ 9.1.1 Length squared importance sampling
The first idea to randomly select columns is to give them importance scores as their norm. This algorithm was introduced by Drineas et al. in [drineas_clustering_2004]. More precisely, the idea is to sample $s$ indices, from a multinomial distribution on $[d]$, where the probability to sample $i$ is $p_{i}\propto\lVert X_{:,i}\rVert^{2}$.
Theorem 3 in [drineas_clustering_2004] states that:
@@ 9.1.2 Leverage score importance sampling
Another importance score that can be considered is the leverage score. This method was introduced by Drineas et al. in [drineas_relative-error_2008]. The $k$-leverage scores distribution is the multinomial distribution on $[d]$ with parameters $p_{i}=l_{i}^{k}/k$. It is easy to see that the leverage scores sum to $k$ so that this definition gives a proper probability distribution. Besides, for $\mathcal{P}_{k}=\operatorname{Span}(V_{k})$, when $(e_{i})_{i\in[d]}$ is the canonical basis of $\mathbb{R}^{d}$, and $\theta_{i}$ is the angle between $\mathcal{P}_{k}$ and $e_{i}$, some linear algebra manipulations show that $l_{i}^{k}=\cos^{2}(\theta_{i})$. So selecting the columns according to the leverage score sampling is in fact choosing the coordinates $e_{i}$ that are the most aligned with $\mathcal{P}_{k}$. PCA selects exactly $\mathcal{P}_{k}$, hence, selecting a subspace which is almost aligned with $\mathcal{P}_{k}$ should yield a good approximation.
Theorem 3 in [drineas_relative-error_2008] states that, for $\epsilon>0$, $\delta>0$, if the number of sampled columns $s$ satisfies $s\geq 4000(k^{2}/\epsilon^{2})\log(1/\delta)$, then under the $k$-leverage score sampling it holds:
@@ 9.2.1 Volume sampling
One way to introduce correlation between the columns of $X$ is through volume sampling. The idea is here to use the kernel $X^{T}X$ and sample a $k$-DPP from it. This method will be called volume sampling (VS) in the following. Notice that the diagonal terms of the kernel are exactly the squared lenghts of the columns of $X$. Volume sampling was first introduced by Deshpande et al. in [deshpande_matrix_2006] and they established the following result:
@@
[[deshpande_matrix_2006] Theorem 1.3] Assume $\mathcal{S}$ is sampled according to the volume sampling method. Then:
@@
Recently, [derezinski_improved_2021] have worked on these bounds and they proved the following result. Notice that the bound is sharp in the worst case scenario, so they play on the decay of the singular values to achieve a better bound in some cases.
Proposition 9.2 ([derezinski_improved_2021] Theorem 1).
@@
Let $0<\epsilon\leq 1/2$, $s<r$ and $t_{s}=s+sr_{s}(X)$, then for $s+\frac{7}{\epsilon^{2}}\ln(\frac{1}{\epsilon^{2}})\leq k\leq t_{s}-1$, the following holds:
@@
This upper bound is rather complicated to interpret in full generality. In their paper [derezinski_improved_2021], the authors show that it actually improves 9.1 in cases where the singular values decay at a polynomial or an exponential rate.
@@ 9.2.2 DPP sampling
The last algorithm we wanted to mention is taken from [belhadji_determinantal_2020]. Their idea is to force the alignment between the subspace generated by the $k$ selected coordinates and the subspace generated by simultaneously force each principle angles to be the smallest possible. In this regard, consider $S\subset[d]$, $\lvert S\rvert=k$ and denote $\mathcal{P}_{S}=\operatorname{Span}\{e_{j}|j\in S\}$. Let $\mathbb{\theta}=(\theta_{i})_{1\leq i\leq k}$ be the principal angle between $\mathcal{P}_{k}$ and $\mathcal{P}_{S}$. Then it can be shown that: $\prod\limits_{i=1}^{k}\cos(\theta_{i})^{2}=\mathrm{Det}(V_{S,[k]}V_{S,[k]}^{T})$. Hence sampling from an elementary DPP (of rank $k$) with kernel $K=V_{k}V_{k}^{T}$ emerges naturally from the geometry of the problem. This is what will be called DPP sampling in the following.
@@
Proposition 9.3 ([belhadji_determinantal_2020] Proposition 16).
@@
Proposition 9.4 ([belhadji_determinantal_2020] Proposition 17).
@@
The proof scheme for this result is quite similar to the one of 9.1 with added technicalities. More precisely, notice that the inequality which was used there is rather loose, and a better understanding of the decay of the singular values of $X$ would give a better bound. Besides, this projection DPP has a strong connection to the principal angles as mentioned previously, and in a certain way, sampling from this DPP makes us find the columns that are best aligned with the principal eigenspaces. Using these facts, and correctly manipulating the symmetric polynomials lead to this result. We refer the reader to [belhadji_determinantal_2020] for more details.
@@
It is worth mentioning that it is possible to reduce the sparsity by looking at some ”effective sparsity”, and conditioning on sampling indices for which the leverage score contributes significantly to the whole. All of these results can also be extended when working with the spectral norm instead of the Frobenius one, see [belhadji_determinantal_2020].
@@ 10.1 Background
Pruning neural networks consists in reducing the number of neurons in the network. Neural network pruning is a central question when it comes to optimizing neural network performances. Mariet and Sra developed a method, DIVNET in [mariet_diversity_2017], for pruning neural network using DPP which has proven to be empirically efficient and is independent of the choice of other parameters such as the activation function or the number of hidden layers. We will describe the method for feed forward neural networks.
@@
Looking at it through the lens of statistical mechanics offers some insight. This framework was already introduced in [acharyya_statistical_2021]. We aim at offering a comprehensive overview of the methods coming from statistical mechanics applied to the study of neural network pruning. We also extend and formalize the results from [acharyya_statistical_2021] to more general settings.
The analysis presented in this section was originally developed in the masters thesis [petrovic2023pruning] of co-author V. Petrovic .
@@ 10.2 The student/teacher framework
The student/teacher framework is a common setting in statistical mechanical analysis of neural networks. The full theory behind this model is fully explained and detailed in [engel_statistical_2001]. The model that we will use in this section comes from [goldt_dynamics_2020] in the noiseless regime.
The framework is the following, assume we are given a data set $(X_{\mu})_{1\leq\mu\leq p}$ of i.i.d. Gaussian variables which take values in $\mathbb{R}^{d}$, of mean 0 and variance matrix $I_{d}$. Let us get two neural networks respectively called the teacher and the student. The teacher network acts as a black box, to which we have access to the activation function and the number of layers and neurons per layer. It gets as inputs the data set $(X_{\mu})_{1\leq\mu\leq p}$ and assigns them outputs $(Y_{\mu})_{1\leq\mu\leq p}$, where if $1\leq\mu\leq p$, $Y_{\mu}=\phi^{*}(X_{\mu})$, where $\phi^{*}$ is a deterministic function. The student network is then another neural network, on which we have full control of the architecture, that we train to reproduce the results of the teacher network. In other words, the student network performs supervised learning with inputs $(X_{\mu},Y_{\mu})_{1\leq\mu\leq p}$. Finally, let $g$ be the activation function, which is chosen to be the sigmoid function $\operatorname{erf}$. For the rest of the section, we will use $i,j,k$ to describe nodes from the student network and $n,m$ to describe nodes from the teacher network.
@@
[…]
@@
The whole analysis will revolve around the dynamics established by Goldt et al. in [goldt_dynamics_2020] .
@@ 10.3 Dynamics of the macroparameters in online learning
Goldt et al establish very general results on the evolution of the paramaters in online learning. In the following, we will focus on a specific case mentioned by the authors of [goldt_dynamics_2019] and further assume that:
@@
[…]
@@
[…]
The authors of [goldt_dynamics_2020] establish the following theorem:
@@ 10.4 Application to neural network pruning
We assume that $T_{nm}=\delta_{nm}$ (as [acharyya_statistical_2021]), i.e., that we work in the ideal case where the weights in the teacher network are a subfamily of orthonormal vectors. Assume further that both hypothesis (H1) and (H2) from section 10.3 hold.
@@
This is a generalization of [acharyya_statistical_2021] in the following sense:
@@ 10.5.3 DPP sampling
Equivalently, if $\in G_{m}, iji\neq j$, then $\mathbb{P}(\{i\in\mathcal{S}\}\cap\{j\in\mathcal{S}\})=0$.
@@
  • the first example is by taking $L_{ij}=Q_{ij}$. Notice that, since $_{ij}=\frac{1}{N}w_{i}^{T}w_{j}, QQ$ is a Gram-matrix so it is semi definite positive, and is an admissible choice for $L$.
@@ 11 A quantum sampler for DPPs
While her thesis initially was concerned with finding a stochastic description of this bunching effect, without the semiclassical description of the electric field as a function-valued stochastic process, Odile Macchi then turned to model a similar situation where photons are replaced by fermions, like electrons. There, she found out that detection events are negatively correlated, evidencing an anti-bunching effect that was later experimentally observed. Macchi formalized DPPs precisely to model that detection process [Mac72]; see also [BFBDHRRSW22Sub] for a modern review. Mathematically, DPPs arise from the combined action of Born’s rule –the quantum physics axiom that lab measurements are draws from a random variable parametrized by two operators on a Hilbert space– and the canonical anti-commutation relations satisfied by the building blocks of operators describing fermions.
In this section, we walk in Macchi’s footprints and illustrate how DPPs are related to quantum fermionic models in a simple setting, showing in passing that we can design an algorithm to sample a given DPP on a quantum computer. This section paraphrases [BaFaFe24], to which we refer the reader for more details and important references, such as [JSKSB18, KePr22].